A full bus holds 24 passengers. Adults pay $25 for the day and children pay $15. On a day when the bus is full the total fees are $540. How many adults and children are there?
Answer by TimothyLamb(4379)

x = adults
y = kids
---
x + y = 24
25x + 15y = 540
---
put the system of linear equations into standard form
---
x + y = 24
25x + 15y = 540
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = adults = 18
y = kids = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---